kiyaR:

How could you use the answers in part a to find powers of i larger than 4, without multiplying them out

TheSmartOne:

Welcome to QuestionCove! What was your answer in part a?

TheSmartOne:

Do you know the power rules? That is what happens to the exponents when you multiply them

kiyaR:

Find i2, i3, and i4. that was my question and my answer i got was -1, -i, 1

TheSmartOne:

Aha, yes of course So really quickly and simply, can you find \(i^5\) \(i^6\) \(i^7\) \(i^8\)

TheSmartOne:

You should notice a very obvious pattern then

kiyaR:

so do you want me to slove what you put up there

TheSmartOne:

Yes please! You can do it faster if you remember because i^4 = 1 so i^5 = i^4 * I And similarly for the other ones

payton:

trust them, they are a genious!! XD

kiyaR:

thanks

payton:

yw

TheSmartOne:

@kiyaR were you able to get it? for example i^6 = i^4 * i^2 And you already know what i, i^2, i^3, i^4 are from your previous question

kiyaR:

yes and I do thank you for helping me get thur it @the smart one you where such a great help

TheSmartOne:

Of course! And in case you need any more clarification, there's a wonderful article that tells you it probably better than I could have on KhanAcademy https://www.khanacademy.org/math/algebra2/x2ec2f6f830c9fb89:complex/x2ec2f6f830c9fb89:imaginary/a/powers-of-the-imaginary-unit

TheSmartOne:

If you have any other questions, feel free to post them and tag me!

kiyaR:

thanks

TheSmartOne:

My pleasure:)

kiyaR:

I will make sure

kiyaR:

In A Midsummer Night’s Dream, the mention of _____ is an allusion to earlier English poetry.

TheSmartOne:

oh my, it's been years since I've read that play. Are there answer choices given?

kiyaR:

yes i will give it to you

kiyaR:

Cupid Demetrius Lysander Titania

TheSmartOne:

mhmm just searched it up and read about it, do you have any idea?

kiyaR:

no not really it gave me an answer and on my odysseyware it say's that it's wrong b/c i chose cupid

TheSmartOne:

Cupid was wrong? oh boi that's what everyone else was saying it was

kiyaR:

i think it's right i'll try again

kiyaR:

thanks anyway

TheSmartOne:

Sure thing!